OpenStudy (anonymous):

Use Stokes' Theorem to find the circulation of F=<2y, 5z, 5x> around the triangle obtained by tracing out the path (4,0,0) to (4,0,4), to (4,4,3) back to (4,0,0).

OpenStudy (anonymous):

Can you define what circulation is?

OpenStudy (anonymous):

i guess it's just going along the path from a point to another?

OpenStudy (anonymous):

I guess they just want you to do: \[ \int_C \mathbf{F}\cdot d\mathbf{r} \]

OpenStudy (anonymous):

i calculated the curl(F) first which was 5, and i multiplied that by 6 which was the region of the triangle but got the wrong answer

OpenStudy (anonymous):

Well I'm guessing circulation is given by: \[ \iint_S \nabla \times \mathbf{F}\cdot d\mathbf{S} \]And by Stokes theorem: \[ \iint_S \nabla \times \mathbf{F}\cdot d\mathbf{S} = \int_C \mathbf{F}\cdot d\mathbf{r} \]

OpenStudy (anonymous):

So basically, just do the line integral. Can you do it?

OpenStudy (anonymous):

but what is the limit of the integral? from 0 to 4?

OpenStudy (anonymous):

You have 3 curves.

OpenStudy (anonymous):

(4,0,0) to (4,0,4) (4,0,4) to (4,4,3) (4,4,3) to (4,0,0)

OpenStudy (anonymous):

Do the line integral across each one and add them up.

OpenStudy (anonymous):

Do you know how to do a line integral?

OpenStudy (anonymous):

Do you know how to parametrize the curves?

OpenStudy (anonymous):

yes! :) i think i get it :P i'm solving it now :) thanks!!!

OpenStudy (anonymous):

Wait... do you suppose it would be easier to find the surface and do the curl...?

OpenStudy (anonymous):

yes but i got the wrong answer

OpenStudy (anonymous):

What was your curl?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Curl is a vector dude....

OpenStudy (anonymous):

curl doted with the normal vector is 5

OpenStudy (anonymous):

that's the integrand

OpenStudy (anonymous):

What was your curl and normal vector?

OpenStudy (anonymous):

5,-5,-2 and 1.0.0

OpenStudy (anonymous):

Also, your surface parametrization was?

OpenStudy (anonymous):

oh i got the right answer! thanks a lot!

OpenStudy (anonymous):

Yeah but I feel like we cheated.